A bullet of mass 55g is shot through a watermelon with mass 1.2kg. the watermelon flies back at 1.5m/s with the bullet inside. what was the bullet’s speed upon entering the watermelon?
1
Expert's answer
2020-02-20T09:24:15-0500

From the conservation of momentum:


"mv=(m+M)u"

"(0.055)v=(0.055+1.2)1.5"

"v=34\frac{m}{s}"